OpenStudy (anonymous):

A dress on sale in a shop is marked at $D. During the discount sale its price is reduced by 15%. Staff are allowed a further 10% reduction on the discounted price. If a staff member buys the dress what will she have to pay in terms of D ?

OpenStudy (anonymous):

Know that the prices are reduced by a percentage of D. So if we get 10% off, it just means that the price we pay is the original D minus the percentage. We can write this as D - 0.10D. D(new) = D(original) - 0.10D with a 10% reduction. In the problem, however, we have two deductions from the original price. One of which is 10%, the other of which is 15%. If you add these deductions you can effectively get 25% off the original price. So: D(new) = D(original) - (0.25D) Now we simplify: D(new) = D - 0.25D D(new) = 0.75D